# Advances in 4D Representation: Geometry, Motion, and Interaction

### v1

@@ 6.2.3 Optimization Objectives
Reconstruction losses optimize photometric consistency using MSE/L1, SSIM, and LPIPS for perceptual sharpness [XLD25, RXM24, LLP25, YCPH25, MCY25]. Extensions include segmentation-aware L1 [LXZ24] and differentiable soft-depth rendering for textured meshes [LLT25]. These losses ensure visual accuracy and serve as universal reconstruction anchors across Gaussian, mesh, and diffusion systems.
@@
Diffusion-based losses minimize noise prediction errors in latent video space [CLZ25, YXV25, JZL25, ZYD25], predicting $\epsilon$ conditioned on frames or camera embeddings. CAT4D [WGP25] uses dual classifier-free guidance for balanced conditioning. AnimateAnyMesh [WYWB25] and DeepVerse [CZH25a] adopt flow-matching objectives for continuous-time supervision. These losses align generated sequences with learned diffusion manifolds, ensuring smooth motion and appearance trajectories.
@@
Geometry losses ensure metric realism. 4DGT [XLD25] uses depth [YKH24] and normal [YQG24] supervision. DrivingRecon [LXZ24] integrates LiDAR depth and 3D position regularization. MonST3R [ZHH24] imposes pointmap alignment and flow-projection consistency, while Geo4D [JZL25] merges multiple geometric terms. Mesh methods [ZYD25, LLT25] use harmonic skinning and dual-quaternion articulation regularization. Interaction methods [LSN24, RTSB25, XLWG23] employ physics-based losses for penetration, contact, and trajectory regularization.
Open challenges. End-to-end training of feed-forward 4D generation frameworks streamline the pipeline by jointly learning geometry, appearance, and motion within a single network. However, this holistic design introduces several limitations: training becomes data-hungry and computationally intensive, optimization tends to be unstable due to intertwined objectives, and fine-grained control over geometry or temporal dynamics is often lost. We discuss some of the drawbacks below:
@@
Training data To train a feedforward model end-to-end there is a requirement of massive scale of 4D data. Explicit 4D data is costly to acquire and less. Most of the existing feed-forward designs [XLD25, RXM24, MCY25, CLZ25] are either trained with a mix of 3D datasets [ZTF18, DLW23] or explicit 4D datasets [GWT24, JTL24] which has limited number of videos. For instance, L4GM [RXM24] relies on 12 M synthetic videos (approx. 300 M frames) and 4DGT [XLD25] on $\sim 5000$ real-world videos. Collecting, curating, and storing such data is costly, while training often spans days on 64–128 GPUs with complex multi-stage optimization. Hence training these methods are infeasible in the academic setup where there are lack of resources to train at this scale.
@@
Limited generalizability Models trained on narrow distributions struggle beyond their training setups—L4GM [RXM24] assumes fixed camera elevation and falters on egocentric or occluded motion; 4DGT [XLD25], tuned for device-specific data, performs poorly on new camera types; and 4DNeX [CLZ25] degrades under severe occlusion or lighting changes. Domain shift from synthetic to real scenes remains an open challenge. For instance, 4D models [RXM24, YXV25, MCY25] trained end-to-end on synthetic datasets like Objaverse-Dy [YXV25] fails to generalize to real-world scenes limiting realism and generalization.
@@
Lack of flexibility Several recent feedforward 4D methods highlight the rigidity of fixed architectural design. [RXM24] directly predicts 3D Gaussian splats from video, tightly coupling the network to a specific representation—changing to meshes or implicit fields would require full retraining. [LRM24] follows a similar feedforward strategy for bullet-time reconstruction, limiting adaptability across representation types. Likewise, 4DGT [XLD25] and St4RTrack [FZW25] demonstrate how feedforward pipelines hard-code 4D Gaussians or pointmap assumptions, reducing flexibility compared to per-scene optimization frameworks that can swap representations or regularizations without retraining.
@@ 6.3 Hybrid Optimization
Two-stage paradigm. A distinct family of 4D methods adopts a two-stage design: Stage 1 performs feed-forward multi-view or video generation, while Stage 2 refines or reconstructs explicit 4D geometry. This separation combines the scalability of diffusion-based appearance synthesis with the precision of geometric optimization. CAT4D [WGP25] first generates view-consistent videos using a multi-view diffusion model (CAT3D + Lumiere), then reconstructs 4D motion via deformable Gaussian splatting. SV4D [XYV24] and SV4D 2.0 [YXV25] similarly use Stable Video Diffusion for multi-view synthesis followed by dynamic NeRF optimization. Splat4D [YCPH25] predicts coarse 3D Gaussians before refining temporal coherence with DynamiCrafter diffusion, while DreamArt [LLT25] synthesizes articulated motion via video diffusion and reconstructs textured meshes through dual-quaternion optimization. Overall, this two-stage feed-forward + optimization framework effectively balances generative realism and geometric consistency, yielding high-fidelity and temporally stable 4D reconstructions.
SDS-Free optimization. Several recent 4D reconstruction and generation methods have deliberately adopted SDS-free training, replacing slow diffusion-based score distillation with direct or self-supervised objectives. Unlike early pipelines such as DreamFusion [PJBM22] and STAG4D [ZJZ24], newer systems employ explicit geometric, photometric, or temporal supervision. Per-scene methods like EG4D [SGW24] and In-2-4D [NCOZMA25] optimize deformation fields using photometric and geometric losses, while AR4D [ZHY25] relies on multi-view reconstruction, flow alignment, and deformation regularization. SV4D [XYV24] and SV4D 2.0 [YXV25] achieve 4D consistency directly within video diffusion frameworks. Data-driven methods such as BTimer [LRM24], 4DGT [XLD25], and DrivingRecon [LXZ24] leverage large-scale video or LiDAR data with depth and flow supervision, making SDS unnecessary. Feedforward Gaussian pipelines—MoVies [LLP25], Splat4D [YCPH25], and Forge4D [HHC25]—use photometric, perceptual, and motion losses for real-time dynamic reconstruction. Overall, these approaches mark a shift from SDS-based optimization toward direct feedforward learning, enhancing scalability, temporal stability, and rendering speed.
@@ 7 Representation Comparison and Trade-off
We compare representations covered so far across six key metrics; see below. Table˜2 outlines the fundamental trade-offs.
@@
Visual Fidelity (Quality of reconstructed appearance, including photorealism and geometric detail preservation). Volumetric rendering (NeRF [PSH21], 3DGS [LWH24]) achieves the highest fidelity through view-dependent effects and continuous radiance modeling. Point clouds [WZH25] and meshes [ZLL25] follow: point clouds directly use input pixel colors despite discrete surfaces, while meshes provide continuous geometry but require PBR modeling for photorealism. Structured representations (graph, part, template [LIC25, LYX24b]) exhibit medium fidelity as appearance depends on node or part attributes rather than direct modeling.
@@ 8 Conclusion, Emerging Trends, and Future Directions
Trend #1: Feed-Forward Reconstruction. The transition from per-scene optimization to feed-forward inference represents a fundamental shift in computational efficiency. Traditional approaches requiring hours of optimization per scene are being replaced by Large Reconstruction Models (LRMs) for 4D through a single forward pass [RXM24, CLZ25], achieving 100-1000× speedups. This trend mirrors similar transitions witnessed in 2D [RBL22] and 3D [ZWZ24a] generation, establishing feed-forward architectures as a dominant paradigm across dimensional barriers.
@@
Trend #3: Integrating World Knowledge. Capturing 4D contents with real-world fidelity requires incorporating high-level semantic understanding and physical plausibility. The use of LLMs has become prevalent for multi-modal reasoning and common-sense guidance in 4D tasks [WYWB25, LIC25], encoding abstract knowledge that would be impractical to specify as mathematical inductive biases or require prohibitive amounts of training data to learn. Complementarily, physics-based constraints are being integrated through differentiable physics losses [LHB25] or reconstruct-then-simulate pipelines [X24, ZYW24], to enforce the generated dynamics to obey fundamental physical laws.

### v2

@@ 6.2.3 Optimization Objectives
Reconstruction losses optimize photometric consistency using MSE/L1, SSIM, and LPIPS for perceptual sharpness [RXM24, MCY25]. Extensions include segmentation-aware L1 [LXZ24] for static-dynamic decoupling. These losses ensure visual accuracy and serve as universal reconstruction anchors across Gaussian, mesh, and diffusion systems.
@@
Diffusion-based losses minimize noise prediction errors in latent video space [CLZ25, JZL25], predicting $\epsilon$ conditioned on frames or camera embeddings. AnimateAnyMesh [WYWB25] adopts flow-matching objectives for continuous-time supervision. These losses align generated sequences with learned diffusion manifolds, ensuring smooth motion and appearance trajectories.
@@
Geometry losses ensure metric realism. 4DGT [XLD25] uses depth [YKH24] and normal [YQG24] supervision. DrivingRecon [LXZ24] employs a DepthNet for depth integration. MonST3R [ZHH24] imposes pointmap alignment and flow-projection consistency, while Geo4D [JZL25] merges multiple geometric terms. Interaction methods [LSN24, RTSB25, XLWG23] employ physics-based losses to penalize interpenetration, enforce contact constraints, and regularize trajectories.
6.2.4 Open Challenges
End-to-end training enables joint learning of geometry, appearance, and motion within a unified framework but introduces limitations in data requirements, generalizability, and architectural flexibility. We discuss some of the drawbacks below:
@@ 6.2.4 Open Challenges
Training data. Feed-forward models require massive-scale 4D data that is costly to acquire. Existing methods [XLD25, RXM24, MCY25, CLZ25] train on mixed 3D datasets [ZTF18, DLW23] or limited explicit 4D datasets [GWT24, JTL24]. L4GM [RXM24] requires 12M synthetic videos (300M frames), while 4DGT [XLD25] uses  5K real-world videos. Training demands days on 64–128 GPUs with complex multi-stage optimization, making academic-scale training infeasible due to computational and storage costs.
@@
Limited generalizability. Models trained on narrow distributions struggle beyond training conditions. L4GM [RXM24] assumes fixed camera elevation and fails on egocentric or occluded motion; 4DGT [XLD25] performs poorly on new camera types; 4DNeX [CLZ25] degrades under occlusion or lighting changes. Domain shift from synthetic to real scenes remains critical—models trained on Objaverse-Dy [YXV25] fail to generalize to real-world scenarios, limiting practical applicability.
@@
Lack of flexibility. Fixed architectural designs limit adaptability. L4GM [RXM24] directly predicts Gaussian splats, tightly coupling the network to this representation—switching to meshes or implicit fields requires full retraining. Similarly, 4DGT [XLD25] and St4RTrack [FZW25] hard-code 4D Gaussians or pointmap assumptions. This contrasts with per-scene optimization frameworks that can swap representations or regularizations without retraining, offering greater experimental flexibility.
@@ 6.3 Hybrid Optimization
Two-stage paradigm. A distinct family of 4D methods adopts a two-stage design: Stage 1 performs feed-forward multi-view or video generation, while Stage 2 refines or reconstructs explicit 4D geometry. This combines the scalability of diffusion-based appearance synthesis with the precision of geometric optimization. CAT4D [WGP25] first generates view-consistent videos using a multi-view diffusion model (CAT3D + Lumiere), then reconstructs 4D motion via deformable Gaussian splatting. SV4D [XYV24] and SV4D 2.0 [YXV25] similarly use Stable Video Diffusion for multi-view synthesis followed by dynamic NeRF optimization. Splat4D [YCPH25] predicts coarse 3D Gaussians before refining temporal coherence with DynamiCrafter, while DreamArt [LLT25] synthesizes articulated motion via video diffusion and reconstructs textured meshes through dual-quaternion optimization. Overall, this two-stage feed-forward + optimization framework balances generative realism and geometric consistency, yielding high-fidelity and temporally stable 4D reconstructions.
SDS-Free optimization. A notable trend in hybrid 4D methods is the shift away from score distillation sampling (SDS) to direct photometric, geometric, and temporal supervision—enabling Stage 2 optimization to remain lightweight and fast. This trend is also broadly observed across per-scene and feed-forward paradigms. Hybrid methods like AR4D [ZHY25], Splat4D [YCPH25], SV4D [XYV24], and SV4D 2.0 [YXV25] rely on flow alignment, uncertainty minimization, and SDS-free NeRF optimization, respectively. Per-scene methods like EG4D [SGW24] and In-2-4D [NCOZMA25] optimize deformation fields using photometric and geometric losses. Feed-forward methods such as BTimer [LRM24], 4DGT [XLD25], DrivingRecon [LXZ24], MoVies [LLP25], and Forge4D [HHC25] leverage depth, flow, photometric, and perceptual losses, making SDS unnecessary.
6.4 Interaction Training
Interaction training aims to generate physically coherent motion sequences for multiple entities from conditioning signals such as text [DD24], audio [SGY24], objects [LWL23], scenes [HCV21], or action classes [XSW23]. Human entities are predominantly modelled with template-based representations—either fully-parametric meshes or skeletons—sometimes augmented or replaced by neural representations such as NeRF and Gaussian Splatting for photorealism, whereas object entities span both structured and unstructured representations. The central challenge lies in capturing not only individual motions but also inter-entity relationships. Existing approaches broadly divide into data-driven and reward-driven paradigms; we briefly overview both below and refer readers to [SGH26] for in-depth discussion.
6.4.1 Data-Driven
Data-driven methods learn to map input conditions to interaction motion sequences from captured data priors (see  Section˜5.3). Architectures have evolved from RNNs [KBM20] and GANs [XSW23] through VAEs [TCBT22, WWZ22] to the now-dominant diffusion models [LZL24, LWL23] and transformers [GZD25]. Training can be further regularized by contact [WZM25], affordance [ZSY25], collision [ZHN20], and physical priors [ZBS24]. While this paradigm produces natural motions grounded in real human data, physical plausibility remains only implicitly enforced, GOAL [TCBT22] and SAGA [WWZ22] resort to post-hoc optimization to reduce penetration, and HOI-Dyn [LW25] introduces explicit driver–responder synchronization via a transformer to improve inter-entity coherence.
6.4.2 Reward-Driven
Reward-driven methods instead train agents via reinforcement learning coupled with a physics simulator, making physical plausibility an intrinsic guarantee rather than a learned prior. The central design problem is the reward function. [HGW23] combines a task-completion reward with a learned realism reward derived from motion data; [WZY25] extends this design to real humanoid robots. SMP [MZS25] repurposes a pretrained motion diffusion model as a frozen, task-agnostic reward, enabling a single prior to guide diverse policies. A key limitation is that most methods train separate per-task policies. [XLWG25] addresses this via teacher–student distillation across diverse HOI datasets, correcting MoCap artifacts through physics simulation. Nevertheless, reward-driven approaches remain constrained by per-task reward engineering and reliance on reference motions, limiting generalization.
6.4.3 Open Challenges
Realism. Template-based body representations cannot capture soft deformation upon contact, limiting realistic grasping and collision-free close interaction. MoCap artifacts such as floating contacts further degrade supervision quality across both paradigms.
Generalization. Neither paradigm generalizes well beyond its training distribution. LLM-guided methods partially address this by decomposing open-vocabulary instructions into structured sub-goals for low-level controllers [XWW23, WPD25, CPH24], yet LLM-generated plans remain physically inconsistent. Tighter integration of LLM semantic reasoning with physics-based execution is a key open direction.
@@ 7 Representation Comparison and Trade-off
We compare representations covered so far across seven key metrics; see below. Table˜2 outlines the fundamental trade-offs.
@@
Visual Fidelity (Quality of reconstructed appearance, including photorealism and geometric detail preservation). Volumetric rendering (NeRF [PSH21], 3DGS [LWH25]) achieves the highest fidelity through view-dependent effects and continuous radiance modeling. Point clouds [WZH25] and meshes [ZLL25] follow: point clouds directly use input pixel colors despite discrete surfaces, while meshes provide continuous geometry but require PBR modeling for photorealism. Structured representations (graph, part, template [LIC25, LYX24b]) exhibit medium fidelity as appearance depends on node or part attributes rather than direct modeling.
@@ 8 Conclusion, Emerging Trends, and Future Directions
Trend #1: Feed-Forward Reconstruction. The transition from per-scene optimization to feed-forward inference represents a fundamental shift in computational efficiency. Traditional approaches requiring hours of optimization per scene are being replaced by Large Reconstruction Models (LRMs) for 4D through a single forward pass [RXM24, CLZ25], achieving 100-1000× speedups. This trend mirrors similar transitions witnessed in 2D [RBL22] and 3D [ZWZ24] generation, establishing feed-forward architectures as a dominant paradigm across dimensional barriers.
@@
Trend #3: Integrating World Knowledge. Capturing 4D contents with real-world fidelity requires incorporating high-level semantic understanding and physical plausibility. The use of LLMs has become prevalent for multi-modal reasoning and common-sense guidance in 4D tasks [WYWB25, LIC25], encoding abstract knowledge that would be impractical to specify as mathematical inductive biases or require prohibitive amounts of training data to learn. Complementarily, physics-based constraints are being integrated through differentiable physics losses [LHB25] or reconstruct-then-simulate pipelines [XZQ24, ZYW24], to enforce the generated dynamics to obey fundamental physical laws.